OpenStudy (anonymous):

11 pictchers, 7 infielders, and 12 other players. if 3 players are selected at ramdom, determine the probl that all 3 are infliders

OpenStudy (anonymous):

OK so we start off the same way!

OpenStudy (anonymous):

Yes, you can do this, it is just like the other two...

OpenStudy (anonymous):

7/30?

OpenStudy (anonymous):

for the first selection, yes, now what for the second? (remember that one infielder is removed by you choosing him.

OpenStudy (anonymous):

6/29

OpenStudy (anonymous):

right, then the third selection...

OpenStudy (anonymous):

5/28

OpenStudy (anonymous):

okay, now multiply...

OpenStudy (anonymous):

so u do 30 times 29 times 28?

OpenStudy (anonymous):

cancel first... it's easier

OpenStudy (anonymous):

then 6 times 5 times 4?

OpenStudy (anonymous):

I mean 7,6 and 5?

OpenStudy (anonymous):

(7*6*5)/(30*29*28) ...the 6*5 cancels with the 30, making it easier

OpenStudy (anonymous):

...=7/(29*28)

OpenStudy (anonymous):

oh, wait keep canceling...

OpenStudy (anonymous):

7/28 = 1/4 so 7/(28*29) = 1/(4*29)

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

7 times 6 times 5 is 210

OpenStudy (anonymous):

yes, but it is easier to cancel before you multiply than to multiply then reduce

OpenStudy (anonymous):

so what would it be when u cancel 210/2?

OpenStudy (anonymous):

do you want to wrangle with reducing 210/24360?

OpenStudy (anonymous):

210/24360 = 1/116

OpenStudy (anonymous):

116?

OpenStudy (anonymous):

Great job on figuring out the probability part, we only got hung up on the arithmetic. When is the quiz?

OpenStudy (anonymous):

Well I am doing h/w now and then next week he usually gives us the quiz like after monday

OpenStudy (anonymous):

then we have a week to work on it